OpenStudy (anonymous):

Factor: 15(x-5)^2-7(x-5)-30

OpenStudy (mertsj):

Let x-5=y 15y^2-7y-30= (3y-5)(5y+6) Now replace y with x-5 [3(x-5)-5][((x-5)+6]= [3x-15-5][x+1) (3x-20)(x+1)

hero (hero):

Luis, I wish you wouldn't use bagatrix to post solutions. You're getting credit for not doing your own work

OpenStudy (mertsj):

Whoops. I left out a 5

hero (hero):

Good solution mertsj

OpenStudy (mertsj):

[(3(x-5)-5][5(x-5)+6]= [3x-15-5][5x-25+6]= (3x-20)(5x-19)

hero (hero):

Well, that's still most likely the best approach to this. It's way better than bagatrix that's for sure

OpenStudy (mertsj):

Thanks Hero.

hero (hero):

Is that a threat? I hope not?

hero (hero):

If a mod saw what you have written it would probably result in an automatic suspension